You want to estimate the proportion of students at OSU who have ever used an online dating
app. You select a random sample of 140 OSU students and find that 43% have used an online
dating app. If you want to construct a 99% confidence interval, what will the margin of error
be? Try not to do a lot of rounding along the way until you get to the end of your calculations.
Answer #1

Solution :

Given that,

n = 140

Point estimate = sample proportion = \hat p =0.43

1 - \hat p = 1-0.43=0.57

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 (\sqrt((0.43*0.57) / 140)

= 0.10778